a resesters of 2 owms plased as a sides of cube and ends of diagnols are connected to a sorce of 15 volts find the eqvanant resester
 

Grade:12

1 Answers

Vikas TU
6 years ago

It can be rearranged as: 
Resistance in first line = r+(1/2r)+(1/2r)+r = 2+1+1+2 = 6 ohm. 
Resistance in second line = (1/2r)+(1/2r) =2 ohm 
Add up to resistance (1/R) = (1/2)+(1/6) = (4/6) 
R = 6/4 = 1.5 ohm.